Patricia deposited $612.40 in a savings account that earns 3.2% simple interest. What is Patricia’s account balance after thirteen years?
Answer by pamalah@yahoo.com(10)

I=P * R * T
I=612.40 * 3.2% * 13
I=$254.76
254.76+612.40=$867.15 in checking account if no other $ deposited but initial deposit and simple interest